OpenStudy (anonymous):

solve the equation. -16 + 2/3 x= -8

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

how did yall get x=12

OpenStudy (anonymous):

Add 16 to both sides Then times by 3 Then divide by 2

OpenStudy (anonymous):

example

OpenStudy (anonymous):

16+2 and 16+3?

OpenStudy (anonymous):

No -16+(2/3)x=-8 (2/3)x=8 2x=24 x=12